OpenStudy (anonymous):

PLEASE HELP LAST PROBLEM! In a region there is a 0.95 probability chance that a randomly selected person of the population has brown eyes. Assume 15 people are randomly selected. a) Find the probability that all of the selected people have brown eyes. b) Find the probability that exactly 14 of the selected people have brown eyes. c) Find the probability that the number of selected people that have brown eyes is 13 or more.

OpenStudy (anonymous):

u got the answer for the last problem.

OpenStudy (anonymous):

yeah thank you for helping!

OpenStudy (anonymous):

for me it came around 9999. 20 something

OpenStudy (anonymous):

hey this problem is same like the before

OpenStudy (anonymous):

i think p(0)=0.05

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

howd you figure that?

OpenStudy (anonymous):

p(1)=15*0.95*0.05^14

OpenStudy (anonymous):

do you know how to find the first one?

OpenStudy (anonymous):

@ganeshie8 @kropot72 this is my last problem then im done for the night

OpenStudy (anonymous):

P = nCx*p^x*q^x use this formula

OpenStudy (anonymous):

see p- probabiltiy of success,and q-probability of failure

OpenStudy (anonymous):

so plug those into that formula above?

OpenStudy (anonymous):

yes .p =95/100 q=5/100

OpenStudy (anonymous):

n=15 so 15c14 *0.95^14*0.05^1

OpenStudy (anonymous):

0.3657 is the answer for case 2)?

OpenStudy (anonymous):

thats for b?

OpenStudy (anonymous):

yes.is it right?

OpenStudy (anonymous):

It wont let me see the answers but i can check another way possibly

OpenStudy (anonymous):

yes! thats right

OpenStudy (anonymous):

for first one a) is 15c15 *0.95^15*0.05^0

OpenStudy (anonymous):

so for first a)=0.4632 right?

OpenStudy (anonymous):

let me see!

OpenStudy (anonymous):

yes!! :)

OpenStudy (anonymous):

then third case can u do that

OpenStudy (anonymous):

thats what im trying to figure out

OpenStudy (anonymous):

the or more might get me but let me take a swing at it

OpenStudy (anonymous):

15c13*0.95^13*0.05^1 ??

OpenStudy (anonymous):

wrong its 15 C13*15c13*0.95^13*0.05^2

OpenStudy (anonymous):

thank you @tejasvir

OpenStudy (anonymous):

anytime:)

OpenStudy (anonymous):

hmm idk if my answer is right?

OpenStudy (anonymous):

i got 14.1489

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

check it again

OpenStudy (anonymous):

i checked it over again and i still have the same answer plugging it in like that @dg2

OpenStudy (anonymous):

i got 0.2695 :-) u should square 0.05

OpenStudy (anonymous):

hmm i must just be plugging it in wrong somehow :/

OpenStudy (anonymous):

but i really appreciate you working me through this whole problem! you helped me out a ton :)

OpenStudy (anonymous):

what was the number after the 5?

OpenStudy (anonymous):

0.269504

OpenStudy (anonymous):

was it?

OpenStudy (anonymous):

okay thanks! i needed to round to 3 decimal places!

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

.964

OpenStudy (anonymous):

but i passed my quiz and thats all that matters :)

OpenStudy (anonymous):

yw:)

OpenStudy (anonymous):

thank you so much!!! :)

OpenStudy (anonymous):

:)